OpenStudy (anonymous):

x2 - 49 = 0 (factoring) Need help not answers

myininaya (myininaya):

x2=49 x=7,x=-7

OpenStudy (anonymous):

how do you get that though? >.<

myininaya (myininaya):

(x-7)(x+7)=0 x=7,x=-7

myininaya (myininaya):

just look at the second way since you said by factoring

OpenStudy (anonymous):

but it doesnt add to -49 that way?

myininaya (myininaya):

if you have a^2-b^2, then the factors are (a-b) and (a+b)

myininaya (myininaya):

what do you mean?

OpenStudy (anonymous):

since the 2 terms are a difference of 2 squares, you can use: \[ (x^2 - y^2) \] \[ (x + y)(x - y) \]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Use that as your guide.

OpenStudy (anonymous):

what if they go like x2 + 24 = 10x?

OpenStudy (anonymous):

Move 10x to the other side and use the quadratic equation : \[ ax^2 + bx + c = 0 \]

OpenStudy (anonymous):

and solve by getting x2 + (add) = (multiply) like as in x2 +16x +48)?

OpenStudy (anonymous):

?

myininaya (myininaya):

x^2+24=10x x^2+24-10x=10x-10x x^2-10x+24=0 (x-6)(x-4)=0

myininaya (myininaya):

if you have x^2+bx+c=0 just find factors of c that add up to be b if you cannot then you cannot factor over the integers

OpenStudy (anonymous):

Then you can use 3 methods in solving for the roots of x. First, by factoring. example: \[ x^2 -4x + 3 = 0 \] \[ (x - 3)(x - 1) = 0 \] Second, by completing the square. example: \[ x^2 -4x = 5\] Find a term that will make the equation a perfect square trinomial the add the term on both sides \[ x^2 -4x +3 = 5 + 3 \] \[ x^2 -4x +3 = 8 \] Third, by using the quadratic equation. \[x = -b \pm \sqrt{b^2 - 4ac}\]------------------ \[ 2a \]

OpenStudy (anonymous):

k now to the next 2x^2 - x - 6

OpenStudy (anonymous):

But i prefer to use factoring because it's the easiest. :)

myininaya (myininaya):

if you have ax^2+bx+c find two factors of a*c that add up to be b so we have 2x^2-x-6 so a=2,b=-1,c=-6 a*c=2(-6)=-4(3) and -4+3=-1(=b) so we replace -x with -4x+3x and then factor by grouping

myininaya (myininaya):

2x^2-4x+3x-6 2x(x-2)+3(x-2) (x-2)(2x+3)

OpenStudy (anonymous):

but the (x)?

OpenStudy (anonymous):

In getting the x, since the starting equation is equated to 0, then equate both factors to 0.

OpenStudy (anonymous):

x - 2 = 0 and 2x + 3 = 0 There you can solve the 2 values of x

OpenStudy (anonymous):

btw before x2 + 24 = 10x? if the middle contains an x? doesnt it show?

OpenStudy (anonymous):

x2 -10x + 24?

OpenStudy (anonymous):

Yes. Using APE, you will have \[ x^2 + 24 -10x = 10x - 10x \] to eliminate x that leads to \[ x^2 + 24 - 10x \] or \[ x^2 -10x + 24 \]

OpenStudy (anonymous):

to move the 10x i mean.

OpenStudy (anonymous):

with myin though its gone?

OpenStudy (anonymous):

x^2+24=10x x^2+24-10x=10x-10x x^2-10x+24=0 (x-6)(x-4)=0 shoudnt their be an x on the last line

OpenStudy (anonymous):

What last line? the very answer? He didn't show the whole solution. he just showed you a part of it for you to work alone.

myininaya (myininaya):

she*

OpenStudy (anonymous):

well give me another example with an x in the middle and show me the steps through it?

OpenStudy (anonymous):

If you're going to continue that' we'll have: x^2+24=10x x^2+24-10x=10x-10x x^2-10x+24=0 (x-6)(x-4)=0 x - 6 = 0 ; x - 4 = 0 x = 6 ; x = 4

OpenStudy (anonymous):

Sorry. lol. :)

OpenStudy (anonymous):

yea i know that but theirs 3 x's

OpenStudy (anonymous):

x^2+24=10x <=== Wheres the ten's x gone?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

the x should not disappear?

OpenStudy (anonymous):

The term with the x will not disapper, it will just be moved to the other side to for the basic quadratic equation ax^2 + bx + c = 0

OpenStudy (anonymous):

disappear.

OpenStudy (anonymous):

x^2-10x+24=0 (x-6)(x-4)=0 the ten's x is not there shouldent it have been like X (x-6) (x-4) or (x2-6) (x-4)?

OpenStudy (anonymous):

im very confused?

OpenStudy (anonymous):

No. It's hard to explain but try to do the FOIL method to check that the factor is right.

myininaya (myininaya):

ok do you believe that -10x=-6x-4x?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok we have x^2-6x-4x+24=0

myininaya (myininaya):

factor by grouping what do the first two terms have in common

myininaya (myininaya):

just look at x^2-6x right now what factors do x^2 and 6x share?

OpenStudy (anonymous):

x?

myininaya (myininaya):

thats right! so we can do x(x-6)-4x+24=0 right?

OpenStudy (anonymous):

yes also did i do 4x2 -10x right with the answer of (2x - 5)?

myininaya (myininaya):

ok what do -4x and 24 have in common?

OpenStudy (anonymous):

That's a different case.

OpenStudy (anonymous):

i dont want to sound stupid but #'s?

myininaya (myininaya):

what is but #s?

OpenStudy (anonymous):

12 ? 4?

OpenStudy (anonymous):

To factor that, you have to get the common factor of each. 4x^2 and -10x have a common factor of 2x so we do it like this : 2x(2x - 5)

myininaya (myininaya):

ok i dont know what you are talking about just go with me for second we have x(x-6)-4x+24=0 this is getting somewhere now what factors do 4x and 24 have in common?

OpenStudy (anonymous):

I have to go now. See you guys.

myininaya (myininaya):

:(

OpenStudy (anonymous):

thank you both

myininaya (myininaya):

hey we aren't done lol what do 4x and 24 have in common?

myininaya (myininaya):

come on derp you can do it

OpenStudy (anonymous):

im done? lol i need help with area now i give up >.>

OpenStudy (anonymous):

(x^2)-49=0 If you have two quantities that are the square of a number being subtracted and equal to 0, then (a+b)(a-b)=(a^2)-(b^2). This general formula will help as you do more problems and get the feel for factoring. It is also helpful to memorize the squares of numbers 1-12 for basic factoring problems. They usually don't get higher than that. In this case, you can say that (a^2)=(x^2), therefore a=x and (b^2)=49. You can find the value of b by taking the square root of each side, to isolate the b from any variables. When you do this, you get b=sqrt49, or b=7. So, put this into the general formula and you get (x+7)(x-7)=0. So, your factors are (x+7) and (x-7) If you need to solve for x, then you set each factor equal to 0 (as stated in the equation) and simply solve it like you would any other algebraic equation.